HUBBLE SPACE TELESCOPE DAILY REPORT #1541 PERIOD COVERED: 0000Z (UTC) 12/20/95 - 0000Z (UTC) 12/21/95 Daily Status Report as of 355/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD 6339 (Verification of the UV Spectrum of the HST = Background Light ) The Faint Object Spectrograph (Red Side) was used to make = ultraviolet observations of the HST background light. These FOS = observations were taken parallel with the Hubble Deep Field (HDF) survey. = Both the HDF and these FOS data will be used to test and improve the = existing engineering model of the HST background light. The observations = were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 0 Successful : 0 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 3 Successful : 3 2.3 Operations Notes: There were no significant opertional problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH